Solve for x and y:
x*2 + y = 7
x + y*2 = 11
I can see that the answer is x = 2, y = 3 BUT how can it be solved other than by graphing?

Answer by stanbon(75887)

Solve for x and y:
x*2 + y = 7
x + y*2 = 11
-----------
Try using substitution:
y=7-x^2
x + (7-x^2)^2=11
x + 49 -14x^2+ x^4 = 11
x^4-14x^2+x+38=0
--------
Using synthetic substitution you can find that
x=2 is the only integer solution.
If x=2, y=3
Cheers,
Stan H.